# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ XII.2.3 Structure Formation
The subsequent analysis will employ Eq. (108) to examine the evolution of density fluctuations. This equation demonstrates that the evolution of a fluctuation $\delta_{\varepsilon}(t,\bm{q})$ with an initial scale $\lambda_{\text{dec}}$, was influenced by three key factors: the expansion of the universe, fluctuations in pressure, and gravitation. Fluctuations in pressure, as indicated by (103), were found to consist of two discrete components: an adiabatic contribution, represented by $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{q})$, and a nonadiabatic contribution, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, where the adiabatic component was initially negligible (115) and the nonadiabatic component was observed to be random.
[…]
Fig. 1 provides a clear summary of the evolution of density fluctuations. It was created using the following methodology. For each value of ${\delta_{T}=-0.01}$, $-0.05$, $-0.10$, $-0.15$, and $-0.20$, Eq. (108) is numerically solved for a large number of values for the initial fluctuation scale $\lambda_{\text{dec}}$ using the initial values (115).
[…]
$d_{\text{H}}(t_{\text{dec}})=a(t_{\text{dec}})\int_{0}^{ct_{\text{dec}}}\!\!\dfrac{\mathrm{d}\tau}{a(\tau)}=3ct_{\text{dec}}\approx 3.5\times 10^{5}\,\text{pc}.$ (117)
This value has been calculated using Eq. (111b) and Tables 1 and 2.
@@
Figure 2: The graphs show the growth rates $\delta^{\prime}_{\varepsilon}$, with initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$, as function of the redshift $z$, or time in million of years. The initial scales $\lambda_{\text{dec}}$ of the fluctuations are measured in parsec. The evolution of density fluctuations started at $z=1090$.
[…]
@@
It is first necessary to determine the particle number density at the time of decoupling denoted by $n_{\scriptscriptstyle(0)}(t_{\text{dec}})$, in order to proceed. This quantity is derived from the Friedmann equation (71a). For $w\ll 1$ and at time $t_{\text{p}}$ this equation is given by:
@@
By employing the relation (111d), one obtains
@@
[…]
@@
Accordingly, the Jeans mass at decoupling is thus given by
@@
As evidenced by the preceding analysis, fluctuations in density, with initial values of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$ and $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$, result in the formation of structures in a flat flrw universe, provided that pressure fluctuations are taken into account. This marks the conclusion of the analysis concerning the evolution of density fluctuations in a flat flrw universe.
XIII Summary and Conclusion
This paper addresses four issues in the field of cosmological perturbation theory for Friedmann-Lemaître-Robertson-Walker (flrw) universes. The first and most significant issue concerns the gauge problem. It is well-established that the linearized Einstein equations and conservation laws yield solutions that are both physical and nonphysical. Consequently, the imposition of physical initial conditions does not result in solutions with physical significance. To resolve this issue, it is crucial to consider gauge-invariant quantities, that is, quantities that are independent of the choice of a coordinate system. It has been demonstrated that perturbations in the energy density and the particle number density can be constructed in a unique manner. In the nonrelativistic limit, it is demonstrated that these quantities represent the actual physical perturbations of the energy density and particle number density in the general theory of relativity. Similarly, as has been done for density perturbations, gauge-invariant perturbations in temperature, pressure, and entropy can be constructed. This addresses the gauge issue in the field of cosmology.
The second issue is that of the growth of minor density fluctuations subsequent to the decoupling of matter and radiation. In a fluid with zero pressure, density fluctuations do not evolve. To address this issue, it is necessary to incorporate the particle number density in addition to the energy density in the linearized Einstein equations and conservation laws. This allows for the kinetic energy density and pressure to be accounted for, resulting in the emergence of local pressure gradients. These gradients give rise to local fluid flows that are instrumental in the formation of structures after decoupling.
A third result is the novel perturbation theory (60), which provides a clear and unambiguous description of the evolution of fluctuations in the particle number density the energy density, as well as their interaction in a perfect fluid where the pressure depends on both the particle number density and the energy density. The aforementioned equations are not only applicable to a flat flrw universe, but also to open and closed universes. Furthermore, the cosmological constant need not be zero. It has been demonstrated that Eqs. (60) yield all previously established physical solutions for large-scale density perturbations in a flat flrw universe in its radiation-dominated phase and in the era following the decoupling of matter and radiation. In the development of the new cosmological perturbation theory, it has been demonstrated that in the nonrelativistic limit, relativistic gauge transformations reduce to Newtonian gauge transformations. Consequently, the conventional relativistic evolution equation—whether derived from general relativity or Newtonian gravitation modified to account for the expansion of the universe—also exhibits gauge-dependent solutions. This finding indicates that Newtonian gravitation is inadequate in accurately describing the evolution of density perturbations.
[…]
@@ XIII Summary and Conclusion
The investigation into cosmological density perturbations in a flat flrw universe reveals that if a nonrelativistic monatomic perfect gas with three degrees of freedom accurately represents the cosmic fluid in our universe following the decoupling of matter and radiation, then a perturbation theory based on general relativity provides an explanation for the origin of the first structures in the universe that does not rely on cdm. In contrast, the theory considers the local nonadiabatic random pressure fluctuations that emerged during the turbulent transition to the era that followed the decoupling of matter and radiation. Subsequent observations with the James Webb Space Telescope are expected to yield novel insights into the conditions that ensued immediately after the decoupling of radiation and matter, as well as the formation of the earliest structures in the universe.
Appendix A Derivation of the Evolution Equations
@@
$\dfrac{p_{\varepsilon}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $\dfrac{p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $H(2-3\beta^{2})$ 00
$\dfrac{\kappa\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}$ 00 $-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $2H+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{6H}$
$\dfrac{-\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ 00 $\dfrac{6\varepsilon_{\scriptscriptstyle(0)}H(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ $\dfrac{\tfrac{3}{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$
In this appendix, the perturbation equations (60) are derived. Prior to undertaking the subsequent analysis, it is necessary to carry out two preliminary steps.
Firstly, as $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$, (4b), it is unnecessary to consider the gauge-dependent quantity $\theta_{{\scriptscriptstyle(1)}}$. As a result, the system (37) can be rewritten in a more suitable form. By eliminating the variable $\theta_{\scriptscriptstyle(1)}$ from the differential equations  (37b)–(37e) using the algebraic constraint equation (37a), a system of four first-order ordinary differential equations is obtained:
@@
Secondly, the energy density and particle number density perturbations (4a) are expressed in the four quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Employing the background equations (10) to eliminate all time derivatives and the linearized constraint equation (37a) to eliminate $\theta_{\scriptscriptstyle(1)}$, the following results are obtained:
@@
From an algebraic standpoint, it is more straightforward to begin by deriving equations for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. This will be accomplished in Sec. A.1. Subsequently, Eqs. (60) for the fluctuations (59) will be derived in Sec. A.2. The expression (44) will be utilized to eliminate the time derivative of $w$.

### v3

@@ XIII.2.3 Structure Formation
The subsequent analysis will use Eq. (108) to examine the evolution of relative density perturbations. This equation shows that the evolution of a relative perturbation, $\delta_{\varepsilon}(t,\bm{q})$, with an initial scale $\lambda_{\text{dec}}$, was influenced by three factors: the expansion of the universe, pressure perturbations, and gravitation. Relative pressure perturbations $\delta_{p}$, as indicated by (103), consisted of two discrete components: an adiabatic contribution, represented by $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{q})$, and a random (102), nonadiabatic contribution, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, where the adiabatic component was initially negligible (115).
Figure 1: The graphs show the redshift and time when a relative perturbation in the energy density with initial scale $\lambda_{\text{dec}}$, and initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$ starting to grow at an initial redshift of $z(t_{\text{dec}})=1090$ has become nonlinear, i.e., when $\delta_{\varepsilon}(t,\bm{q})=1$. The graphs are labeled with the initial values of the random, nonadiabatic pressure perturbations $\delta_{T}(t_{\text{dec}},\bm{q})$. For each graph, the Jeans scale is $6.4\,\text{pc}$.
Since $w\ll 1$, (92), the nonadiabatic component could have taken on relatively large positive or negative values for small differences in the relative perturbations of energy and particle number density, as shown in (101). Consequently, there were regions where the total relative pressure perturbation was initially negative after decoupling. This allowed density perturbations to grow rapidly, despite the fact that the expansion of the universe opposed growth. As the adiabatic component of the relative pressure perturbation increased, the total relative pressure perturbation became positive. This significantly reduced the growth rate of density perturbations. The initial growth phase was brief yet long enough for a density perturbation to reach the nonlinear regime several hundred million years after the Big Bang. Conversely, if the nonadiabatic relative pressure perturbation was positive initially, then the total relative pressure perturbation was also positive. This resulted in the formation of voids, which were regions with negative values of $\delta_{\varepsilon}$. In these regions, matter was driven to the edges.
Fig. 1 provides a summary of the evolution of relative density perturbations. It was created using the following methodology. For each value of $\delta_{T}=-0.01$}, 0.05-0.05, 0.10-0.10, 0.15-0.15, and 0.20-0.20, Eq. (108) was solved numerically for a large number of values for the initial perturbation scale $\lambda_{\text{dec}}$ using the initial values (115). The integration started at $z=1090$ and ended when either $z=0$ or when $\delta_{\varepsilon}=1$. Each integration run returned a single point on the graph for a particular choice of the initial scale $\lambda_{\text{dec}}$ when $z>0$ and $\delta_{\varepsilon}=1$. In this case, it is clear that the perturbation has become nonlinear within $13.79\,\text{Gyr}$. On the other hand, if the integration stopped at $z=0$ and $\delta_{\varepsilon}<1$, then the perturbation had not yet reached its nonlinear phase today. Each graph shows the time and scale at which $\delta_{\varepsilon}=1$ for a given value of $\delta_{T}$.
As illustrated in Fig. 1, the optimal scale for growth was approximately $6.4\,\text{pc}$. Perturbations at this scale entered the nonlinear phase as early as $40$ million years ($z\approx 49$) after the Big Bang. Perturbations with scales smaller than $6.4\,\text{pc}$ reached their nonlinear phase at a much later time because their internal gravitation was weaker than for large-scale perturbations. Furthermore, pressure perturbations and expansion of the universe led to oscillatory behavior, as shown in Fig. 2. In contrast, perturbations with a scale greater than $6.4\,\text{pc}$ were less affected by pressure perturbations. However, because of their larger size, the expansion worked against their growth, so that, despite their stronger gravitation, they also reached the nonlinear phase at a later time. Perturbations larger than $70\,\text{pc}$ did not reach their nonlinear regime within $13.79\,\text{Gyr}$.
Perturbations that became nonlinear within $13.79\,\text{Gyr}$ are within the particle horizon at decoupling given by:
$d_{\text{H}}(t_{\text{dec}})=a(t_{\text{dec}})\int_{0}^{ct_{\text{dec}}}\!\!\dfrac{\mathrm{d}\tau}{a(\tau)}=3ct_{\text{dec}}\approx 3.5\times 10^{5}\,\text{pc},$ (116)
where Eq. (111b) and Tables 1 and 2 have been used.
@@
Figure 2: The graphs show the growth rates $\delta^{\prime}_{\varepsilon}$, with initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$, as function of the redshift $z$, or time in million of years. The initial scales $\lambda_{\text{dec}}$ of the perturbations are measured in parsec. The evolution of relative density perturbations started at $z=1090$.
As Fig. 2 shows, the largest initial growth rates occurred for perturbations that were smaller than $6.4\,\text{pc}$. The smallest perturbations which had become nonlinear within $13.79\,\text{Gyr}$, had a scale of approximately $2\,\text{pc}$, as shown in Fig. 1. Perturbations with scales $2\,{\text{pc}}\lesssim\lambda_{\text{dec}}\lesssim 4.1\,\text{pc}$} oscillated toward the nonlinear phase within $13.79\,\text{Gyr}$. After approximately $14$ million years, the total pressure perturbation (103) had become positive and the growth rate had decreased, resulting in gravitation and, to a lesser extent, expansion becoming the primary drivers of the evolution of density perturbations. Consequently, the most turbulent phase of density perturbations had concluded, and a phase of gradual and consistent gravitational growth toward the nonlinear regime had commenced. As can be seen in Fig. 1, at a redshift of $z\approx 7$, or $600$ million years after the Big Bang, all perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$ had become nonlinear.
Assuming a spherical density perturbation with a diameter equal to $\lambda_{\text{dec}}$, the mass of the perturbation at decoupling is given by
@@
The particle number density at the time of decoupling, $n_{\scriptscriptstyle(0)}(t_{\text{dec}})$, can be derived from the Friedmann equation (75a). For $w\ll 1$ and at time $t_{\text{p}}$ this equation is given by:
@@
By using the relation (111d), one obtains
@@
The relations (117) and (119) and the constants in Tables 1 and 2 reveal that perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$, had masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$.
Because of the steepness of the graphs in Fig. 1 for scales below $6.4\,\text{pc}$, this scale is designated as the relativistic Jeans scale. Its value is
@@
Accordingly, the Jeans mass at decoupling is given by
@@
The above analysis shows that, if nonadiabatic pressure perturbations are taken into account, relative density perturbations with initial values $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$ and $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$} can give rise to structure in the early universe after the decoupling of matter and radiation.
XIV Summary and Conclusion
This article introduces a relativistic cosmological perturbation theory (62) that describes the evolution of relative energy density and particle number density perturbations in closed ($K=+1$), flat ($K=0$), and open (K=1{K=-1}) Friedmann-Lemaître-Robertson-Walker (flrw) universes, where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=-6K/a^{2}$ is the spatial curvature. It consists of two differential equations: a second-order inhomogeneous evolution equation for the relative energy density perturbation and a first-order entropy equation that determines the evolution of the source term of the first equation. These equations together describe the evolution of nonadiabatic relative perturbations $\delta_{\varepsilon}(t,\bm{x}):=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})/\varepsilon_{{\scriptscriptstyle(0)}}(t)$ and $\delta_{n}(t,\bm{x}):=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})/n_{{\scriptscriptstyle(0)}}(t)$ in energy and particle number densities, respectively.
The entropy equation implies that before matter and radiation decoupled, both ordinary matter perturbations and Cold Dark Matter (cdm) perturbations are coupled to energy density perturbations by gravitation. Consequently, cdm did not initiate structure formation after decoupling.
After decoupling, pressure gradients became essential for fluid flow, which is crucial for forming structures, since in a pressureless fluid, density perturbations $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ are static.
Newtonian gravity is not a suitable framework for studying the evolution of cosmological density perturbations. This is because the conventional evolution equation, whether derived from general relativity or Newtonian gravity, yields gauge-dependent solutions. Consequently, no distinction exists between sub- and super-horizon density perturbations.
XIV.1 Results for a Flat FLRW Universe
In a radiation-dominated universe, relative density perturbations $\delta_{\varepsilon}$ oscillated with an amplitude that increased proportionally to the square root of time.
After matter and radiation decoupled at $t=t_{\text{dec}}$, the cosmological fluid became a nonrelativistic perfect fluid. According to observations made by the Planck satellite [50], relative density perturbations were of the order of $10^{-5}$ at the time of decoupling. These perturbations are far too small to explain structures in the universe by gravitational growth, if only adiabatic perturbations are considered. This is because the growth of an adiabatic density perturbation slows down due to the increasing pressure perturbation caused by the density perturbation itself. For this reason, this article examines nonadiabatic density perturbations. This was achieved using an equation of state for a nonrelativistic perfect gas that considers not only the particles’ rest energy, but also their kinetic energy. Therefore, in addition to the energy density, the particle number density must also be considered. Then, the total pressure is given by $\delta_{p}(t,\bm{x})\approx\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})+\delta_{T}(t_{\text{dec}},\bm{x})$, where $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ is the adiabatic component caused by the density perturbation itself, and a random nonadiabatic component $\delta_{T}(t_{\text{dec}})$. In addition, the solution to the entropy equation is identical to the perturbed equation of state of the energy density, expressed in terms of relative density perturbations and the nonadiabatic component of pressure, that is, $\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})\approx-\tfrac{3}{2}w(t)\delta_{T}(t_{\text{dec}},\bm{x})$, where $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}\ll 1$. This expression shows that small random fluctuations in both $\delta_{n}(t_{\text{dec}},\bm{x})$ and $\delta_{\varepsilon}(t_{\text{dec}},\bm{x})$ — arising from the rapid and chaotic transition to the post-decoupling era — could lead to significant local random fluctuations in the relative nonadiabatic pressure perturbation, $\delta_{T}(t_{\text{dec}},\bm{x})$, which could take on both positive and negative values. If the total relative pressure perturbation $\delta_{p}(t_{\text{dec}},\bm{x})$ was positive at the moment of decoupling, a void was formed and matter moved toward its edges. Conversely, if $\delta_{p}(t_{\text{dec}},\bm{x})$ was negative, the density perturbation would undergo a brief period of rapid growth. This rapid growth continued until the total relative pressure perturbation $\delta_{p}(t,\bm{x})$ became positive. After that, the density perturbation slowly but steadily grew toward the nonlinear phase, which was reached early on. Perturbations smaller than $70\,\text{pc}$ would reach their nonlinear phase within $13.79$ billion years or less. The relativistic Jeans scale was $6.4\,\text{pc}$, and the corresponding Jeans mass was $2.2\times 10^{4}\,\text{M}_{\odot}$. Perturbations of Jeans scale size reached their nonlinear regime at a redshift of $z\approx 49$ or approximately $40$ million years after the Big Bang. All density perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$ reached their nonlinear phase within $z\approx 7$ or $600$ million years. These scales correspond to masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$. This is smaller than what was observed by the James Webb Space Telescope (jwst) [51], which observed mature galaxies. Nevertheless, it is reasonable to assume that the masses continued to increase in the nonlinear phase due to their greater gravitational pull.
XIV.2 Conclusion
@@
Investigating cosmological density perturbations in a flat flrw universe reveals that if a nonrelativistic, monatomic, perfect gas with three degrees of freedom accurately represents the cosmological fluid in our universe after matter and radiation decoupled, then a perturbation theory based on general relativity can explain the origin of the first structures in the universe without relying on cdm. Subsequent observations with the jwst, launched on Christmas Day $2021$, are expected to yield novel insights into the conditions that ensued immediately after the decoupling of radiation and matter, as well as the formation of the earliest structures in the universe. These observations will help us determine if general relativity accurately describes the formation of the first structures in the universe.
Appendix A Derivation of the
Perturbation Theory
@@ Appendix A Derivation of the
$\dfrac{p_{\varepsilon}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $\dfrac{p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $H(2-3\beta^{2})$ 0
$\dfrac{\kappa\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}$ 0 $2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $2H+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{6H}$
$\dfrac{-\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ 0 $\dfrac{6\varepsilon_{\scriptscriptstyle(0)}H(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ $\dfrac{\tfrac{3}{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$
In this appendix, the perturbation theory (62) is derived. Before undertaking this task, it is necessary to carry out two preliminary steps.
Firstly, as $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$, (6b), it is unnecessary to consider the gauge-dependent quantity $\theta_{{\scriptscriptstyle(1)}}$. Eliminating the quantity $\theta_{{\scriptscriptstyle(1)}}$ from the differential equations (37b)–(37e) using the algebraic constraint equation (37a) yields a system of four first-order ordinary differential equations suitable for computer algebra software:
@@
Secondly, the energy density and particle number density perturbations (6a) are expressed in the four quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Using the background equations (11) to eliminate all time derivatives, and the linearized constraint equation (37a) to eliminate $\theta_{\scriptscriptstyle(1)}$, the following results are obtained:
@@
From an algebraic standpoint, it is more straightforward to begin by deriving equations for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. This will be accomplished in Sec. A.1. Subsequently, Eqs. (62) for the relative perturbations (61) will be derived in Sec. A.2. The expression (44) will be used to eliminate the time derivative of $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$.